ts 2req DRs viow Topics Use the References to access important values if needed for this question. Complete the table below for calculating the molar mass of the compound tetraphosphorus decaoxide. Molar mass Moles Mass in one mole of tetraphosphorus decaoxide g/mol mol g/mol mol Molar mass tetraphosphorus decaoxide- g/mol Submit Answer Retry Entire Group 9 more group attempts remaining
Answer #1

Molar mass of P4O10,

MM = 4*MM(P) + 10*MM(O)

= 4*30.97 + 10*16.0

= 123.88 + 160

= 283.88 g/mol

Complete the table below for calculating the molar mass of the compound tetraphosphorus decaoxid Molar mass Moles Mass in one